# Training-and-prompt-free General Painterly Harmonization Using Image-wise Attention Sharing

### v1

@@ 6. Conclusion
In this work, we introduce a novel share-attention layer, enabling the utilization of image-wise attention from different images. Furthermore, we devised the similarity reweighting technique capable of flexibly controlling the attention strength of reference images without the need for fine-tuning or prompt. Based on these innovations, we propose TF-GPH able to perform a more general form of the painterly harmonization task including object insertion, object swapping and style transfer. Furthermore, we construct the GPH Benchmark along with range-based evaluation aiming to mitigate the current shortage of evaluations for image editing methods. We believe that TF-GPH provides ordinary users with more convenience for content creation, and our contributions can facilitate growth in future research on image editing methods. In the future, we plan to further explore the potential extensions of share-attention layer to other attention-based applications and modalities, such as 3D generation (Mildenhall et al., 2021; Cai et al., 2023; Lin et al., 2023) and video creation (Ho et al., 2022b, a; Blattmann et al., 2023).
@@ Appendix A Appendix: Algorithm
The TF-GPH framework is based on stable-diffusion (SD) (Rombach et al., 2022), combined with DPM-solver (Lu et al., 2022), which not only reduces the timestep requirements but also supports the functionality of the inversion process. And this inversion process can be further stablized with a fixed prompt embedding $\rho_{\text{except}}$. We assume that the inversion process for the TF-GPH input has been completed. In the subsequent reconstruction stage, cross-image information is incorporated into the output image via our share-attention module. Our proposed share-attention module is a plug-and-play component, designed to replace the attention layer in the original SD framework. Furthermore, Our focus lies in the share-attention layer’s forward function (share-FORWARD), where we only substitute the forward mechanism of the original attention layer while retaining trained parameters (Q-K-V Projection layer, normalization layer, etc.). The detailed algorithm for TF-GPH is outlined in Alg. 1.
Data: initial noise $Z_{T}=cancat\{z_{T}^{\text{ref1}},z_{T}^{\text{ref2}},z_{T}^{\text{comp}}\}$, step $T_{\text{share}}$ and layer depth $L_{\text{share}}$ to start using share attention module, the similarity weight $\hat{\alpha}$ and $\hat{\beta}$
@@ 15       $O_{l+1}\leftarrow$ $O_{l}$+$CA_{l}$($O_{l}$,C) ;
Appendix B Appendix: Datasets
B.1. WikiArt (Tan et al., 2019) with COCO (Lin et al., 2014).
This dataset has been widely adopted by various general stylization methods (Lu et al., 2023a; Cao et al., 2023a; Huang et al., 2023; Deng et al., 2022) due to its high flexibility and feasibility. Following common evaluation practices, we utilize the WikiArt dataset for background images and the COCO dataset as the source for foreground objects. Specifically, we randomly sampled 1000 images from the WikiArt validation dataset and 1000 segmented objects across 80 different classes from the COCO validation dataset (with each object’s class equally distributed). These segmented objects are then composited onto the background images to generate our final composite images. (The evaluation result can be found in Fig. 16,Tab. 4).However, a limitation of this dataset lies in its diversity, which is constrained to combinations of real-world objects (from COCO) against paintings (typically European style). As a result, the data distribution from WikiArt combined with COCO may not fully represent real-world applications of image harmonization tasks, which often involve combining fictional objects with various forms of graphic art.
B.2. TF-ICON Benchmark (Lu et al., 2023b).
@@ Appendix B Appendix: Datasets
B.3. General Painterly Harmonization Benchmark.
@@
Figure 7. The demo of GPH-Benchmark, consisting of more general test data falling out of distribution of previous training-required methods (Lu et al., 2023b, a; Meng et al., 2022). First row: ”insert object”, ”background”, ”Ours”. Second row: ”TF-ICON” (Lu et al., 2023b),”PHDiff” (Lu et al., 2023a), ”SDEdit” (Meng et al., 2022)
Appendix C Appendix: Baselines
@@
PHDiffusion (Lu et al., 2023a): PHDiffusion (abbreviated as PHDiff in subsequent sections) is a framework for painterly harmonization. They propose incorporating an additional adaptive encoder combined with a fusion module into the existing stable diffusion pipeline and fine-tuning this combined pipeline on the WikiArt with COCO dataset. It is noteworthy that within this framework, they introduce an additional ”Strength” hyperparameter to control the scale of the fusion module, which also represents the influence of the background style on the pasted object. Consequently, to evaluate the performance of the content emphasis strategy of PHDiff, we set the ’Strength’ parameter to 0, while for the stylization emphasis strategy, we set it to 1. (Their default setting for ’Strength’ is fixed at 0.7).
@@
TF-ICON (Lu et al., 2023b): The TF-ICON approach primarily leverages two hyperparameters to govern the prompt-based image composition process. Firstly, $\tau_{\alpha}$ indicates the onset of attention injection (0 for beginning, 1 for the end), and secondly, $\tau_{\beta}$ denotes the onset of the rectification process. Unfortunately, neither $\tau_{\alpha}$ nor $\tau_{\beta}$ directly modulates stylization intensity. However, reducing $\tau_{\alpha}$ generally yields more stylized outputs, while reducing $\tau_{\beta}$ produces images closer to the composite. Thus, to assess TF-ICON’s content emphasis strategy, we follow the setting suggest by TF-ICON, we set $\tau_{\alpha}=0.4$ and $\tau_{\beta}=0$ as proposed in their paper for photography composition, demanding heightened content preservation. Conversely, for the stylization emphasis strategy, we adopt $\tau_{\alpha}=0.4$ and $\tau_{\beta}=0.8$ as recommended for cross-domain composition, necessitating higher stylization strength.
VCT (Cheng et al., 2023): The VCT method utilizes a set of specially designed loss to fine-tune the given stable-diffusion model with one-shot style reference. Consequently, the longer the tuning process takes, the more the SD model fits to the given style. Thus, to assess VCT’s content emphasis strategy, we directly discard the tuning process using default SD model for reconstruction, as for the stylization emphasis strategy, we simply leverage their default hyperparameters for tuning, which take around 5 to 10 minutes tuning process for a single RTX 3090 GPU on each test sample (while the inference time for single image usually take 5 to 30 seconds).
Appendix D Appendix: More Qualitative Result
D.1. Qualitative of WikiArt w/ COCO
We provide the comparison of our proposed method toward other baselines in Fig. 16.
D.2. Qualitative of TF-Benchmark
We provide the comparison of our proposed method toward other baselines in Fig. 8.
@@ D.2. Qualitative of TF-Benchmark
Figure 8. Results on TF-Benchmark. Our results (second row) show better content preservation than TF-ICON (first row).
D.3. More Qualitative result of GPH Benchmark
We provide more result of TF-GPH on GPH Benchmark in Fig. 17 (Insertion and swapping) and Fig. 18 (Style Transfer).
Appendix E Appendix: More Quantitative Result
E.1. Quantitaive of WikiArt w/ COCO
@@
We provide the comparison of range-based evaluation on WikiArt combined with COCO in Tab. 4.
Appendix F Appendix: User Study
@@
Survey flow. At the outset, each participant will receive instructions and a demonstration question aimed at familiarizing them with the answer flow and evaluation criteria (”Content Consistency,” ”Style Similarity,” ”Visual Quality”). Subsequently, they will be required to answer 10 randomly selected questions from a pool of 30 questions, along with an attention-check question designed to assess the validity of their responses (details provided in the following section). Furthermore, the options for each question will be shuffled for enhanced reliability. The participant’s view of the question is illustrated in Fig. 9.
@@
Figure 9. A painterly harmonization question participant might need to answer.
Attention-check question. The intent behind these questions is to confirm the participant’s grasp of the concepts of ”content consistency” and ”style similarity.” For instance, within the construction of attention-check questions for the style transfer task, we provide the content reference and style reference directly, expecting them to be ranked at first place for ”content consistency” and ”style transfer,” respectively, by the participant. As illustrated in Fig. 10, where the participant should rank (C) in the first place of content consistency, and (D) in the first place of style similarity. We reject the responses that fails the attention check and collect 50 valid responses for each painterly harmonization and style transfer task.
Figure 10. The attention-check question for style transfer task.
Appendix G Appendix: More Experiment
G.1. Experiment of Inference-time-adjustable Hyperparameters
@@
The TF-GPH incorporates four adjustable hyperparameters during inference: $T_{\text{share}}$, $T_{\text{L}}$, $\alpha$, and $\beta$. The initial two parameters, $T_{\text{share}}$ and $T_{\text{L}}$, govern the commencement timing of the share-attention layer; an earlier start of the share-attention layer leads to increased blending of objects into the background. Meanwhile, the latter two parameters, $\alpha$ and $\beta$, regulate the weighting of references in the reconstruction process; decreasing $\alpha$ and increasing $\beta$ result in a more stylized output. These adjustments afford TF-GPH enhanced adaptability across a diverse array of usage scenarios. We present a straightforward visualization depicting the shift in quantitative indices solely by varying $T_{\text{share}}$ in Fig. 11, alongside the qualitative outcome of altering others Fig. 12 and Fig. 13.
@@
Figure 11. Comparison of different stylized strength, when adjusting $T_{share}$ only.
@@
Figure 12. Comparing various levels of stylized strength by adjusting both $T_{\text{share}}$ and $L_{\text{share}}$ (abbreviated as ”S”), with fixed values for $\alpha=0.9$ and $\beta=1.1$.
@@
Figure 13. Comparing various levels of stylized strength by adjusting both $\alpha$ and $\beta$, with fixed values for $T_{\text{share}}=15$ and $L_{\text{share}}=7$.
Appendix H Appendix: More Applications
H.1. Inpainting

### v2

@@ Conclusion
In this work, we introduce a novel similarity disentangle mask, faciliatating the utilization of attention from different images. Furthermore, we devised the similarity reweighting technique capable of controlling the attention strength of reference images without the need for fine-tuning or prompt. Based on them, we propose TF-GPH to perform a more general form of painterly harmonization. Also, we construct the GPH Benchmark with range-based evaluation aiming to mitigate the current shortage of evaluations for image editing. Both human and quantitative evaluations show that TF-GPH produces more harmonious results, which should benefit future research in image editing.
Acknowledgments
This work is partially supported by the National Science and Technology Council, Taiwan under Grants NSTC-112-2221-E-A49-059-MY3 and NSTC-112-2221-E-A49-094-MY3.
@@ Appendix A Appendix: Algorithm
The TF-GPH framework is based on stable-diffusion (SD) (Rombach et al. 2022), combined with DPM-solver (Lu et al. 2022), which not only reduces the timestep requirements but also supports the functionality of the inversion process. And this inversion process can be further stablized with a fixed prompt embedding $\rho_{\text{except}}$. We assume that the inversion process for the TF-GPH input has been completed. In the subsequent reconstruction stage, cross-image information is incorporated into the output image via our share-attention module. Our proposed share-attention module is a plug-and-play component, designed to replace the attention layer in the original SD framework. Furthermore, Our focus lies in the share-attention layer’s forward function (share-FORWARD) with additional similarity disentangle mask $\hat{M}$, where we only substitute the forward mechanism of the original attention layer while retaining trained parameters (Q-K-V Projection layer, normalization layer, etc.). The detailed algorithm for TF-GPH is outlined in Alg. 1.
Data: initial noise $Z_{T}=cancat\{z_{T}^{\text{f}},z_{T}^{\text{b}},z_{T}^{\text{c}}\}$, step $T_{\text{share}}$ and layer depth $L_{\text{share}}$ to start using share attention module, the similarity weight $\hat{\alpha}$ and $\hat{\beta}$
@@ 15       $O_{l+1}\leftarrow$ $O_{l}$+$CA_{l}$($O_{l}$,C) ;
Appendix B Appendix: Visualization
We visualize how the similarity reweighting technique change the attention of $z^{\text{c}}$ toward two different sources $z^{\text{f}}$ and $z^{\text{b}}$. We perform the visualization of attention on layer 14 of UNet in Fig.7 below. With the similarity disentangle mask only (columns 1, 2), the information from foreground dominate the attention of inserted object, when the similarity reweighting is included (columns 3, 4), the background information significantly influences the inserted object from the early denoising step t=15 till the last step t=0.
Figure 7: The visualization of how simialrty reweighting change the attention.
We also visualize the impact of our similarity reweighting technique on the feature similarity distribution between the composite image and the content and style references, as illustrated in Fig. 8. Our first observation is that similarity reweighting effectively reduces the influence of the content reference, thereby creating more space for stylization. Furthermore, we validate our claim that reweighting the similarity before the softmax operation significantly enhances the influence of high-similarity features while diminishing that of low-similarity features. As shown in the right part of Fig. 8, although style features predominantly exhibit low similarity values (less than 0.1), the increase in similarity is relatively greater for high-similarity features (values above 0.2).
Figure 8: The visualization of how simialrty reweighting change the attention distribution.
Appendix C Appendix: Datasets
WikiArt (Tan et al. 2019) with COCO (Lin et al. 2014).
This dataset has been widely adopted by various general stylization methods (Lu et al. 2023; Cao, Hong, and Niu 2023; Huang et al. 2023; Deng et al. 2022) due to its high flexibility and feasibility. Following common evaluation practices, we utilize the WikiArt dataset for background images and the COCO dataset as the source for foreground objects. Specifically, we randomly sampled 1000 images from the WikiArt validation dataset and 1000 segmented objects across 80 different classes from the COCO validation dataset (with each object’s class equally distributed). These segmented objects are then composited onto the background images to generate our final composite images. (The evaluation result can be found in Fig. 16, Tab. 7 and Tab. 8).However, a limitation of this dataset lies in its diversity, which is constrained to combinations of real-world objects (from COCO) against paintings (typically European style). As a result, the data distribution from WikiArt combined with COCO may not fully represent real-world applications of image harmonization tasks, which often involve combining fictional objects with various forms of graphic art.
TF-ICON Benchmark (Lu, Liu, and Kong 2023).
@@ Appendix C Appendix: Datasets
General Painterly Harmonization Benchmark.
@@
Appendix D Appendix: Baselines
ProPIH (Niu et al. 2024b): ProPIH is a novel painterly harmonization pipeline, different from previous autuencoder-based methods, they design a multi-stage harmonization network, which harmonize the composition foreground from low-level style to high-level style. We directly choose the first stage output as the stylization lower bound, while the last stage output as stylization output.
@@ Appendix D Appendix: Baselines
PHDiffusion (Lu et al. 2023): PHDiffusion (abbreviated as PHDiff in subsequent sections) is a framework for painterly harmonization. They propose incorporating an additional adaptive encoder combined with a fusion module into the existing stable diffusion pipeline and fine-tuning this combined pipeline on the WikiArt with COCO dataset. It is noteworthy that within this framework, they introduce an additional ”Strength” hyperparameter to control the scale of the fusion module, which also represents the influence of the background style on the pasted object. Consequently, to evaluate the performance of the content emphasis strategy of PHDiff, we set the ’Strength’ parameter to 0, while for the stylization emphasis strategy, we set it to 1. (Their default setting for ’Strength’ is fixed at 0.7).
@@
TF-ICON (Lu, Liu, and Kong 2023): The TF-ICON approach primarily leverages two hyperparameters to govern the prompt-based image composition process. Firstly, $\tau_{\alpha}$ indicates the onset of attention injection (0 for beginning, 1 for the end), and secondly, $\tau_{\beta}$ denotes the onset of the rectification process. Unfortunately, neither $\tau_{\alpha}$ nor $\tau_{\beta}$ directly modulates stylization intensity. However, reducing $\tau_{\alpha}$ generally yields more stylized outputs, while reducing $\tau_{\beta}$ produces images closer to the composite. Thus, to assess TF-ICON’s content emphasis strategy, we follow the setting suggest by TF-ICON, we set $\tau_{\alpha}=0.4$ and $\tau_{\beta}=0$ as proposed in their paper for photography composition, demanding heightened content preservation. Conversely, for the stylization emphasis strategy, we adopt $\tau_{\alpha}=0.4$ and $\tau_{\beta}=0.8$ as recommended for cross-domain composition, necessitating higher stylization strength.
ZSTAR (Deng et al. 2023): They reveal that the cross-attention mechanism in latent diffusion models tends to blend the content and style images, resulting in stylized outputs that deviate from the original content image. To overcome this issue, they introduce a cross-attention rearrangement strategy, for stlization lowe bound we restrict this rearrangement only to the middle 16th attention layer, as for stylization upper bound, we allow these arrangement in all the 1st to 32th attention layer.
StyleID (Lu et al. 2023): Furthermore they introduce query preservation and attention temperature scaling to mitigate the issue of disruption of original content and initial latent Adaptive Instance Normalization (AdaIN) to deal with the disharmonious color, they already provide the default settings, for stylization lower bound they recommend setting gamma to 0.75. As for stylization upper bound, they suggest the setting of gamma to 0.3.
Appendix E Appendix: More Qualitative Result
Qualitative of WikiArt w/ COCO
We compare our proposed method with common WikiArt w/ COCO baselines in Fig. 16. As shown, our method produces convincing outputs across a wide range of styles, including combinations of objects such as humans, food, and animals merged with styles like Impressionism, Modern sketch, and Abstract painting. Unlike other baselines that primarily focus on matching the color tones of objects with the background, our method better utilizes the existing elements of the background reference, such as brushstrokes, abstract edges, and inherent colors. This results in more coherent and harmonious outputs, outperforming existing models.
Qualitative of TF-Benchmark
We provide the comparison of our proposed method toward other baselines in Fig. 9. Although TF-ICON produces harmonious outputs, the content often becomes distorted. For instance, the panda’s pose is changed, the hamburger’s content is altered, and the tower’s shape is modified. In contrast, our method not only better preserves the identity of the objects but also seamlessly blends them into the reference background. Showing the advantage of TF-GPH in the aspect of content identity preservation.
@@ Appendix E Appendix: More Qualitative Result
Figure 9: Results on TF-Benchmark. Our results (second row) show better content preservation than TF-ICON (first row).
More Qualitative result of GPH Benchmark
We provide more result of TF-GPH on GPH Benchmark in Fig. 17 (Insertion and swapping) and Fig. 18 (Style Transfer). These example contains novel objects which are not included in the common COCO dataset, such as pyramid, cartton character, and seal. We also include the sumi-e, cartoon and menga background reference serving as content or style references as shown in the last 3 row in Fig. 18 (Style Transfer). These examples validate the efficacy of TF-GPH methods upon uncommon input, which are often out of the training data of common painterly harmonization dataset.
Appendix F Appendix: More Quantitative Result
Quantitaive of GPH Benchmark in 256x256
Our proposed TF-GPH is a novel pipeline designed for generating images at resolutions of 512x512 and higher. In contrast, existing methods like ArtoPIH and ProPIH are limited to generating images at 256x256 resolution. To thoroughly evaluate the performance of these models in low-resolution scenarios, we resize the output of our methods to facilitate painterly harmonization at lower resolutions Tab. 6. As shown in the table, our TF-GPH method still produce competative result especially in the stylization related metrics.
Quantitaive of WikiArt w/ COCO
@@ Appendix F Appendix: More Quantitative Result
We provide the comparison of range-based evaluation on WikiArt combined with COCO in Tab. 7 (512x512) and Tab. 8 (256x256) .
Quantitative Evaluation on TF-ICON Benchmark
We also evaluate the performance of our proposed TF-GPH on the existing prompt-guided image composition benchmark, TF-ICON. As shown in Tab. 3 , our method achieves state-of-the-art performance on this benchmark, even without the support of prompts.
Table 3: Quantitative result of TF-ICON Benchmark
SDEdit (0.4) 0.35 0.62 80.56 27.73
Blended 0.11 0.77 73.25 25.19
Paint 0.13 0.73 80.26 25.92
DIB 0.11 0.63 77.57 26.84
TF-ICON 0.10 0.60 82.86 28.11
Ours 0.05 0.48 83.34 30.33
Appendix G Appendix: User Study
@@
Survey flow. At the outset, each participant will receive instructions and a demonstration question aimed at familiarizing them with the answer flow and evaluation criteria (”Content Consistency,” ”Style Similarity,” ”Visual Quality”). Subsequently, they will be required to answer 20 randomly selected questions from a pool of 60 questions, along with an attention-check question designed to assess the validity of their responses (details provided in the following section). Furthermore, the options for each question will be shuffled for enhanced reliability. The participant’s view of the question is illustrated in Fig. 10.
@@
Figure 10: A painterly harmonization question participant might need to answer.
Appendix H Appendix: Inference-time-adjustable Hyperparameters
Experement Seeting
For the style transfer task, we set $\alpha$ to 0.9, $\beta$ to 1.1, $T_{share}$ to 25 (we activate the share-attention layer when t¡$T_{share}$), and $L_{share}$ to 14 (out of 16 total layers in the diffusion U-Net). As for the object swapping purpose, we change $T_{share}$ to 20, and for the object insertion usage, we change $T_{share}$ to 15. We refer to these hyperparameters ($T_{share}$, $L_{share}$, $\alpha$, $\beta$) as “inference-time-adjustable hyperparameters” since they can be flexibly adjusted to modulate the strength of style according to different use cases during the inference process. As we have emphasized, there is no universally optimal harmonization sweet point for all aesthetic preferences; it depends entirely on the specific use case. The only quantitative metrics we can establish are the lower and upper bounds of stylization. Therefore, instead of exhaustively evaluating every possible combination of hyperparameters, we selected settings that produce relatively harmonious outputs for our preference.
Sensitive Test
@@ Appendix H Appendix: Inference-time-adjustable Hyperparameters
The TF-GPH incorporates four adjustable hyperparameters during inference: $T_{\text{share}}$, $T_{\text{L}}$, $\alpha$, and $\beta$. The initial two parameters, $T_{\text{share}}$ and $T_{\text{L}}$, govern the commencement timing of the share-attention layer; an earlier start of the share-attention layer leads to increased blending of objects into the background. Meanwhile, the latter two parameters, $\alpha$ and $\beta$, regulate the weighting of references in the reconstruction process; decreasing $\alpha$ and increasing $\beta$ result in a more stylized output. These adjustments afford TF-GPH enhanced adaptability across a diverse array of usage scenarios.
Why $\gamma=-\infty$
We set $\gamma=-\infty$ to simplify content and style disentanglement. As shown in the table Tab.4, using other values like 0.9, 1, or 1.1 lower due to the introduction of additional content-related features. With $\gamma=-\infty$, content and style are directly controlled by $\alpha$ and $\beta$, as supported by Figs.4(a), (b) and ”+SDM” in Tab.2, where reconstruction difference are negligible.
Effect of $(\alpha,\beta)$
The quantitative results are shown Tab.5, with qualitative examples shown in Fig. 13 and our project page. We observe that decreasing $\alpha$ or $\beta$ leads to grayish tones, while increasing them results in over-saturated colors with minimal stylization gains. This is reflected quantitatively: other settings disrupt background consistency, increasing with minor improvement. Thus, we chose $(\alpha,\beta)=(0.9,1.1)$ to maintain background consistency while enhancing object stylization, which are the goal of painterly harmonization. Also we present a straightforward visualization depicting the difference by varying $T_{share}$ as shown in Fig. 13 $T_{\text{share}}$ in Fig. 11, alongside the qualitative outcome of altering others Fig. 12 and .
$(\alpha,\beta,\gamma)$ $LP_{bg}$ $LP_{fg}$ $CP_{style}$
$(1,1,-\infty)$ 0.11 0.10 47.50
$(0.9,1.1,-\infty)$ 0.12 0.32 56.37
$(1,1,1)$ 0.11 0.10 47.50
$(0.9,1.1,0.9)$ 0.12 0.25 52.33
$(0.9,1.1,1)$ 0.11 0.13 48.71
$(0.9,1.1,1.1)$ 0.11 0.10 47.61
Table 4: Ablation study on the additional entry $\gamma$. The results indicate that incorporating $\gamma$ does not further expand the lower or upper bounds.
$(\alpha,\beta)$ (1,1) (0.9, 1.1) (0.9,1.5) (0.5,1.5) (1.5, 1.5) (0.5, 0.5)
$LP_{bg}$ 0.11 0.12 0.18 0.19 0.22 0.26
$LP_{fg}$ 0.10 0.32 0.36 0.37 0.14 0.27
$CP_{style}$ 47.50 56.37 55.96 56.48 45.64 49.51
Table 5: Ablation study on different settings of stylization strength. We observed that increasing the difference between $(\alpha,\beta)$ marginally enhances stylization strength but introduces significantly more noise, compromising background preservation.
@@
Figure 11: Comparison of different stylized strength, when adjusting $T_{share}$ only.
@@
Figure 12: Comparing various levels of stylized strength by adjusting both $T_{\text{share}}$ and $L_{\text{share}}$ (abbreviated as ”S”), with fixed values for $\alpha=0.9$ and $\beta=1.1$.
@@
Figure 13: Comparing various levels of stylized strength by adjusting both $\alpha$ and $\beta$, with fixed values for $T_{\text{share}}=15$ and $L_{\text{share}}=7$.
Appendix I Appendix: More Applications
These examples emerged as part of our broader exploration into the full potential of our proposed TF-GPH method. As our design has the property of harmoniously mixing the features from two different images by scaling the similarity (rather than directly adjusting attention output), we can perform tasks such as semantic mixing and exemplar-based inpainting. These toy examples are included in the Appendix to share results that might inspire similar lines of research and also as part of our future work.
Inpainting